OpenStudy (anonymous):

-1/12 - 3/4 = A. -1/3 B. -5/6 C. 2/3

OpenStudy (anonymous):

b

OpenStudy (ash2326):

we have \[-\frac{1}{12}-\frac{3}{4}\] multiply numerator and denominator of second term by 3 \[-\frac{1}{12}-\frac{3\times 3}{4\times 3}\] we get \[-\frac{1}{12}-\frac{9}{12}\] now subtract the two terms we get \[-\frac{10}{12}\] DIVIDE BOTH NUMERATOR AND DENOMINATOR BY 2 WE GET \[-\frac{5}{6}\]